# Attention as an Adaptive Filter

### v2

@@ 3.7 Normalized Queries & Keys
If the query-key and value dynamics parameters are shared, $\mathbf{P}^{C}_{\mathbf{qk}}=\mathbf{P}^{C}_{\mathbf{v}}:=\mathbf{P}^{C}$, so:
$\mathbf{\tilde{A}}=\frac{\nu\mathbf{I}}{\nu\mathbf{V}^{C}+\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}}$
@@
$(\nu\mathbf{V}^{C}+\big|\mathbf{R}_{\mathbf{qk}}\big|^{2})[i,j]=\nu\big|\lambda_{C}\big|^{2}\sigma^{2}\frac{1-e^{-2\alpha\Delta t_{ij}}}{-2\alpha}+\eta^{2}e^{-2\alpha\Delta t_{ij}}+1+e^{-2\alpha\Delta t_{ij}}$
@@
Collecting terms and letting $a=1+\frac{\nu|\lambda_{C}|^{2}\sigma^{2}}{-2\alpha}$, $b=1+\eta^{2}-\frac{\nu|\lambda_{C}|^{2}\sigma^{2}}{-2\alpha}$, this can be written as:
$(\nu\mathbf{V}^{C}+\big|\mathbf{R}_{\mathbf{qk}}\big|^{2})[i,j]=a+be^{-2\alpha\Delta t_{ij}}-2e^{-\alpha\Delta t_{ij}}\sum_{k}\mathrm{Re}\left(\mathbf{\tilde{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\tilde{Z}}_{\mathbf{k}}[k,j]\right)$
@@
3.8 Unitary Dynamics
@@
$\sigma_{V}^{2}(\tau):=\big|\lambda_{C}\big|^{2}\sigma^{2}\Delta t\tau+\eta^{2}$
@@
$\log(\nu\mathbf{V}^{C}+2)-\log\left(1-\frac{2\mathrm{Re}\big(\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\big)}{\nu\mathbf{V}^{C}+2}\right)$
@@
$\mathbf{L}\approx-\log(\nu\mathbf{V}^{C}+2)+\frac{2\mathrm{Re}\big(\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\big)}{\nu\mathbf{V}^{C}+2}$
If we set $\sigma^{2}\rightarrow 0$ (zero process noise), i.e. assume a deterministic system, then $\mathbf{V}^{C}=\eta^{2}\mathbf{1}$ becomes a simple scalar bias. Letting $\tau=2/(\nu\eta^{2}+2)$, the unnormalized softmax becomes:
@@
The constant term $\frac{\tau}{2}$ outside the exponential cancels in the normalization. The imaginary part of the matrix exponential becomes a multiplicative relative positional encoding, of the form of xPos (Sun et al., 2022) and RetNet (Sun et al., 2023). Under the assumption $\alpha=0$, the propagated covariance $\mathbf{V}^{C}$ is a bias of the form a|ij|+ba|i-j|+b, as in ALiBi (Press et al., 2022).
In the above formulation, if the inputs are normalized beforehand, then the projection matrices $\mathbf{W_{q}}$ and $\mathbf{W_{k}}$ must be unitary to ensure that $|\mathbf{\hat{Z}}_{q}|$ and $|\mathbf{\hat{Z}}_{k}|$ remain 1, simplifying the computation. By contrast, normalizing the queries and keys after the projections removes this constraint, since $\mathbf{W_{q}}$ and $\mathbf{W_{k}}$ no longer need to preserve norms. This motivates the choice of post-projection normalization, a strategy explored in prior work such as Query-Key Normalization for Transformers (Henry et al., 2020).
3.9 The Inhomogeneous Case
@@
3.10 Measurement Time Delay and Bias
@@ 4.1 Algorithm
We use scalar noise parameters $\sigma^{2},\eta^{2}$ and dynamics parameters $\alpha,\bm{\omega}$, where the real part of the eigenvalues is set to $\alpha^{2}$ to ensure stability of our LTI model. Note that the eigenvalues of $\bm{\Omega}$ appear in complex conjugate pairs, as a necessary condition for $\mathbf{A}$ to be real, so we can write the vector of eigenvalues as:
@@
We switch from the rational weight $\tfrac{1}{1+r_{ij}^{2}/\nu}$ to the equivalent exponential–log form $\exp\!\big(-\tau_{s}\log(1+r_{ij}^{2}/\nu)\big)$, where $\tau_{s}$ is the usual temperature parameter that controls the sharpness of the weighting, with the aim of improving numerical stability; this reduces exactly to the rational form when $\tau_{s}=1$. We drop $\big|\lambda_{C}\big|^{2}$ since we can absorb it into $\sigma^{2}$, and drop the Mahalanobis scale $\nu$, since we can absorb it into $\sigma,\eta,\tau_{s},\mathbf{W}_{Q},\mathbf{W}_{K}$.
Note that $\bm{\Omega}$ is similar to a block diagonal matrix of $2\times 2$ rotation matrices, i.e. there exists some $\mathbf{P}\in\mathbb{R}^{d\times d}$ such that:
[…]
However, to retain our fully diagonalized representation, we must work with complex numbers. We define the following linear layer variant to reproduce the effect of complex multiplication:
Complex Linear Layer. We denote a complex linear layer by $\mathcal{L}(\cdot)$. Instead of operating directly on complex vectors, we represent inputs and outputs by stacking their real and imaginary parts. For an input $\mathbf{x}=\left[\mathbf{x}_{r},\mathbf{x}_{i}\right]^{\top}$ with $\mathbf{x}_{r},\mathbf{x}_{i}\in\mathbb{R}^{d}$, the layer computes:
$\mathcal{L}(\mathbf{x})=\begin{bmatrix}\mathbf{W}_{r}&-\mathbf{W}_{i}\\[4.0pt] \mathbf{W}_{i}&\mathbf{W}_{r}\end{bmatrix}\mathbf{x}+\begin{bmatrix}\mathbf{b}_{r}\\[4.0pt] \mathbf{b}_{i}\end{bmatrix}$,
where $\mathbf{W}_{r},\mathbf{W}_{i}\in\mathbb{R}^{d_{e}\times d}$ are the real and imaginary weight matrices, and $\mathbf{b}_{r},\mathbf{b}_{i}\in\mathbb{R}^{d_{e}}$ are real-valued biases (where $d_{e}$ is the embedding dimension). This construction is equivalent to multiplication by a complex weight matrix $\mathbf{W}_{r}+i\mathbf{W}_{i}$ with bias $\mathbf{b}_{r}+i\mathbf{b}_{i}$, while remaining entirely in the real domain.
$\mathbf{W}_{ij}:=\mathbf{W}_{r,ij}+i\mathbf{W}_{i,ij}=\alpha_{ij}\begin{bmatrix}\cos(\phi_{ij})\\ \sin(\phi_{ij})\end{bmatrix},\quad\alpha_{ij}\sim\sqrt{\tfrac{2}{d_{\text{in}}+d_{\text{{out}}}}}\,\mathcal{N}(0,1),\;\;\phi_{ij}\sim\mathcal{U}(0,2\pi).$
Since the inputs to the network layer are real-valued, we can concatenate a vector of zeros to the inputs to represent the imaginary part. In general, the outputs of the network layer may be complex-valued. We pass the real part of the output to the next layer and discard the imaginary part. Alternatively, we can design the input and output projections to map between real and complex inputs/outputs. We pursue the former option here, for simplicity.
Another option to enforce real inputs is to impose a penalty on $\big|\operatorname{Im}(\mathbf{Z_{p}})\big|$ or on $\big|\mathbf{W_{v}}\mathbf{W_{p}}-\mathbf{I}^{d_{v}\times d_{v}}\big|$. For simplicity, we do not impose penalties and allow these matrices to freely vary.
@@
In our implementation, queries and keys are computed from the current values rather than directly from the raw inputs. This design enables a natural recursive approximation of full batch attention: as the value estimates are iteratively updated at each step, the queries and keys reflect the latest state, allowing each iteration to propagate residual information through the trajectory. This is optional.
@@
For simplicity, we use shared dynamic parameters $\alpha,\bm{\omega}$ for the queries/keys and values, so $\bm{\tilde{\Phi}}_{\mathbf{qk}}^{+}=\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}$ and $\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}=\bm{\tilde{\Phi}}_{\mathbf{v}}^{-}$. We define these as separate variables however, since in general they can be different.
Input: Input sequence $\mathbf{Z}_{r}\in\mathbb{R}^{d\times m}$
Complex linear layers: $\mathcal{L}_{Q},\mathcal{L}_{K},\mathcal{L}_{V},\mathcal{L}_{P}$; dynamics/noise parameters: $\alpha,\bm{\omega},\sigma,\eta$; time step $\Delta t$; softmax temperature $\tau_{s}$; residual weight $\delta$.
@@
Lift real inputs to complex by concatenating zeros: $\mathbf{Z}=\left[\mathbf{Z}_{r},\mathbf{0}\right]^{\top}\in\mathbb{C}^{2\times d\times m}$
$\big(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}})\big)\leftarrow\mathcal{L}_{V}\big(\operatorname{Re}(\mathbf{Z}),\operatorname{Im}(\mathbf{Z})\big).$
$\big(\operatorname{Re}(\mathbf{Z_{q}}),\operatorname{Im}(\mathbf{Z_{q}})\big)\leftarrow\mathcal{L}_{Q}\big(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}})\big)$,
$\big(\operatorname{Re}(\mathbf{Z_{k}}),\operatorname{Im}(\mathbf{Z_{k}})\big)\leftarrow\mathcal{L}_{K}\big(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}})\big)$.
@@
Covariance kernel: $\sigma^{2}_{V}(\tau)=\sigma^{2}\kappa[|\tau|]+\eta^{2}\mathcal{K}^{2}_{\alpha}[|\tau|],\quad\mathbf{V}^{C}[i,j]=\sigma^{2}_{V}(i-j)$
@@
$\mathbf{L}=-\log\left(\mathbf{V}^{C}+\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}+\epsilon\right)\odot\mathbf{M}_{\text{causal}},$
@@
Prediction: $\mathbf{Z}_{\mathbf{v}}^{+}=(e^{\alpha\Delta t}e^{\bm{\lambda}_{\Omega}\Delta t})\odot\mathbf{\bar{Z}_{v}}$
Output Projection: $\big(\operatorname{Re}(\mathbf{Z_{p}}),\operatorname{Im}(\mathbf{Z_{p}})\big)\leftarrow\mathcal{L}_{P}\big(\operatorname{Re}(\mathbf{Z}_{\mathbf{v}}^{+}),\operatorname{Im}(\mathbf{Z}_{\mathbf{v}}^{+})\big)$
Return: $\operatorname{Re}(\mathbf{Z_{p}})$ ;
// We discard $\operatorname{Im}(\mathbf{Z_{p}})$.
@@
    Complex-valued multiplicative positional encodings with learned parameters;
@@
    Scalar decay constants (one per head) with pair-wise uncertainty propagation via the closed-form solution of the differential Lyapunov equation;

### v3

@@ 3.7 Further Simplification: Normalized Queries & Keys
If the query-key and value dynamics parameters are shared, $\mathbf{P}^{Z}_{\mathbf{qk}}=\mathbf{P}^{Z}_{\mathbf{v}}:=\mathbf{P}^{Z}$, so:
$\mathbf{\tilde{A}}=\frac{\nu\mathbf{I}}{\nu\mathbf{V}^{Z}+\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}}$
@@
$(\nu\mathbf{V}^{Z}+\big|\mathbf{R}_{\mathbf{qk}}\big|^{2})[i,j]=\nu\big|\lambda_{C}\big|^{2}\sigma^{2}\frac{1-e^{-2\alpha\Delta t_{ij}}}{-2\alpha}+\eta^{2}e^{-2\alpha\Delta t_{ij}}+\gamma^{2}+1+e^{-2\alpha\Delta t_{ij}}$
@@
Collecting terms and letting $a=1+\gamma^{2}+\frac{\nu|\lambda_{C}|^{2}\sigma^{2}}{-2\alpha}$, $b=1+\eta^{2}-\frac{\nu|\lambda_{C}|^{2}\sigma^{2}}{-2\alpha}$, this can be written as:
$(\nu\mathbf{V}^{Z}+\big|\mathbf{R}_{\mathbf{qk}}\big|^{2})=a\mathbf{I}+b\mathbf{E}_{\mathbf{qk}}^{2}-2\mathbf{E}_{\mathbf{qk}}\odot\mathrm{Re}\big(\mathbf{\tilde{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\tilde{Z}}_{\mathbf{k}}\big)$
@@
3.8 Further Simplification: Unitary Dynamics
@@
$\sigma_{V}^{2}(\tau):=\big|\lambda_{C}\big|^{2}\sigma^{2}\Delta t\tau+(\eta^{2}+\gamma^{2})$
@@
$\log(\nu\mathbf{V}^{Z}+2)-\log\left(1-\frac{2\mathrm{Re}\big(\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\big)}{\nu\mathbf{V}^{Z}+2}\right)$
@@
$\mathbf{L}\approx-\log(\nu\mathbf{V}^{Z}+2)+\frac{2\mathrm{Re}\big(\mathbf{\hat{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\big)}{\nu\mathbf{V}^{Z}+2}$
If we set $\sigma^{2}\rightarrow 0$ (zero process noise), i.e. assume a deterministic system, then $\mathbf{V}^{Z}=(\eta^{2}+\gamma^{2})\mathbf{1}$ becomes a simple scalar bias. Letting $\tau=2/(\nu(\eta^{2}+\gamma^{2})+2)$, the unnormalized softmax becomes:
@@
The constant term $\frac{\tau}{2}$ outside the exponential cancels in the normalization. We recover a complex-valued generalization of dot-product attention. The imaginary part of the state transition matrix $\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}$ functions as a generalized rotary-positional encoding (RoPE) (Su et al., 2023). We discuss this connection in Section 6 (Related Work).
In the above formulation, if the inputs are normalized beforehand, then the projection matrices $\mathbf{W_{q}}$ and $\mathbf{W_{k}}$ must be unitary to ensure that $|\mathbf{\hat{Z}}_{\mathbf{q}}|$ and $|\mathbf{\hat{Z}}_{\mathbf{k}}|$ remain 1, simplifying the computation. By contrast, normalizing the queries and keys after the projections removes this constraint, since $\mathbf{W_{q}}$ and $\mathbf{W_{k}}$ no longer need to preserve norms. This motivates the choice of post-projection normalization, a subject we will return to later.
3.9 Extension: Inhomogeneous Dynamics
@@
3.10 Extension: Measurement Time Delay and Bias
@@ 4 Implementation
4.1 Complex-valued Computations
To map real-valued inputs $\mathbf{Z}\in\mathbb{R}^{d\times m}$ to complex-valued queries and keys, we can define linear layers $\mathcal{L}_{\mathbb{R}\rightarrow\mathbb{C}}(\mathbf{Z}):\mathbb{R}^{d\times m}\rightarrow\mathbb{R}^{2d\times m}$, as follows:
$\mathcal{L}^{\mathbb{R}\rightarrow\mathbb{C}}(\mathbf{Z})=\begin{bmatrix}\mathbf{W}_{r}\\[4.0pt] \mathbf{W}_{i}\end{bmatrix}\mathbf{Z}+\begin{bmatrix}\mathbf{b}_{r}\\ \mathbf{b}_{i}\end{bmatrix}$,
where $\mathbf{W}_{r},\mathbf{W}_{i}\in\mathbb{R}^{d\times d}$, $\mathbf{b}_{r},\mathbf{b}_{i}\in\mathbb{R}^{d}$.
The $d$-dimensional complex-valued queries, keys, and values are then represented as $2d$-dimensional real-valued matrices:
$\mathbf{Z}_{\mathbf{q}}=\mathcal{L}_{\mathbf{q}}^{\mathbb{R}\rightarrow\mathbb{C}}(\mathbf{Z}),\quad\mathbf{Z}_{\mathbf{k}}=\mathcal{L}_{\mathbf{k}}^{\mathbb{R}\rightarrow\mathbb{C}}(\mathbf{Z}),\quad\mathbf{Z}_{\mathbf{v}}=\mathcal{L}_{\mathbf{v}}^{\mathbb{R}\rightarrow\mathbb{C}}(\mathbf{Z}),$
$\mathbf{Z_{q}}=\begin{bmatrix}\mathbf{Z}_{\mathbf{q},r}\\ \mathbf{Z}_{\mathbf{q},i}\end{bmatrix},\quad\mathbf{Z_{k}}=\begin{bmatrix}\mathbf{Z}_{\mathbf{k},r}\\ \mathbf{Z}_{\mathbf{k},i}\end{bmatrix},\quad\mathbf{Z_{v}}=\begin{bmatrix}\mathbf{Z}_{\mathbf{v},r}\\ \mathbf{Z}_{\mathbf{v},i}\end{bmatrix}\,\,\in\mathbb{R}^{2d\times m}$
Define cosine and sine matrices:
$\mathbf{C}_{\mathbf{qk}}[k,i]=\cos(\omega_{\mathbf{qk},k}t_{i}),\quad\mathbf{S}_{\mathbf{qk}}[k,i]=\sin(\omega_{\mathbf{qk},k}t_{i})$
Then, the complex rotations are given by:
$\mathbf{\tilde{Z}_{q}}=\begin{bmatrix}\mathbf{\tilde{Z}}_{\mathbf{q},r}\\ \mathbf{\tilde{Z}}_{\mathbf{q},i}\end{bmatrix}=\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}\odot\mathbf{Z_{q}}=\begin{bmatrix}\mathbf{C}_{\mathbf{qk}}\odot\mathbf{Z}_{\mathbf{q},r}+\mathbf{S}_{\mathbf{qk}}\odot\mathbf{Z}_{\mathbf{q},i}\\ \mathbf{C}_{\mathbf{qk}}\odot\mathbf{Z}_{\mathbf{q},i}-\mathbf{S}_{\mathbf{qk}}\odot\mathbf{Z}_{\mathbf{q},r}\end{bmatrix}$
for the queries (and similarly for the keys and values).
We compute the complex inner products as:
$\mathrm{Re}(\mathbf{\tilde{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\tilde{Z}}_{\mathbf{k}})=\begin{bmatrix}\mathbf{\tilde{Z}}_{\mathbf{q},r}\mathbf{\tilde{Z}}_{\mathbf{k},r}+\mathbf{\tilde{Z}}_{\mathbf{q},i}\mathbf{\tilde{Z}}_{\mathbf{k},i}\end{bmatrix}$
And the moduli of the queries and keys as:
$|\mathbf{Z}_{\mathbf{q}}|^{2}=\mathbf{Z}_{\mathbf{q},r}^{2}+\mathbf{Z}_{\mathbf{q},i}^{2},\quad|\mathbf{Z}_{\mathbf{k}}|^{2}=\mathbf{Z}_{\mathbf{k},r}^{2}+\mathbf{Z}_{\mathbf{k},i}^{2}$
Since the attention matrix $\mathbf{\hat{A}}$ is real-valued, we have:
$\mathbf{\hat{A}}\mathbf{\tilde{Z}}_{\mathbf{v}}^{\dagger}=\begin{bmatrix}\mathbf{\hat{A}}\mathbf{\tilde{Z}}_{\mathbf{v},r}\\ -\mathbf{\hat{A}}\mathbf{\tilde{Z}}_{\mathbf{v},i}\end{bmatrix}:=\begin{bmatrix}\mathbf{Z}_{\mathbf{a},r}\\ \mathbf{Z}_{\mathbf{a},i}\end{bmatrix}$
Then, the estimate is:
$\mathbf{\bar{Z}}_{\mathbf{v}}^{\dagger}=\begin{bmatrix}\mathbf{\tilde{Z}}_{\mathbf{v},r}\\ -\mathbf{\tilde{Z}}_{\mathbf{v},i}\end{bmatrix}=(\bm{\tilde{\Phi}}_{\mathbf{v}}^{+})^{\dagger}\odot(\mathbf{\hat{A}}\mathbf{\tilde{Z}}_{\mathbf{v}}^{\dagger})=\begin{bmatrix}\mathbf{C}_{\mathbf{v}}\odot\mathbf{Z}_{\mathbf{a},r}-\mathbf{S}_{\mathbf{v}}\odot\mathbf{Z}_{\mathbf{a},i}\\ \mathbf{C}_{\mathbf{v}}\odot\mathbf{Z}_{\mathbf{a},i}+\mathbf{S}_{\mathbf{v}}\odot\mathbf{Z}_{\mathbf{a},r}\end{bmatrix}$
The prediction is computed in the same way. Define single time-step cosine and sine vectors:
$\mathbf{c}_{\Delta t}[k]=\cos(\omega_{k}\Delta t),\quad\mathbf{s_{v}}[k]=\sin(\omega_{k}\Delta t)$
And broadcast these to be of dimension $d\times m$:
$\mathbf{C}_{\Delta t}=\mathbf{c}_{\Delta t}\mathbf{1}_{1\times m},\quad\mathbf{S}_{\Delta t}=\mathbf{s}_{\Delta t}\mathbf{1}_{1\times m}$
The prediction is then:
$\mathbf{Z}_{\mathbf{v}}^{+}=e^{\alpha\Delta t}e^{\bm{\omega}\Delta t}\odot\mathbf{\bar{Z}_{v}}=e^{\alpha\Delta t}\begin{bmatrix}\mathbf{C}_{\Delta t}\odot\mathbf{Z}_{\mathbf{a},r}+\mathbf{S}_{\Delta t}\odot\mathbf{Z}_{\mathbf{a},i}\\ \mathbf{C}_{\Delta t}\odot\mathbf{Z}_{\mathbf{a},i}-\mathbf{S}_{\Delta t}\odot\mathbf{Z}_{\mathbf{a},r}\end{bmatrix}$
To map the complex-valued $\mathbf{Z}_{\mathbf{v}}^{+}$ back into the original basis, we can use another linear layer $\mathcal{L}_{\mathbb{C}\rightarrow\mathbb{R}}:\mathbb{R}^{2d\times m}\rightarrow\mathbb{R}^{d\times m}$, as follows:
$\mathcal{L}^{\mathbb{C}\rightarrow\mathbb{R}}(\mathbf{Z}_{\mathbf{v}}^{+})=\begin{bmatrix}\mathbf{W}_{r}&-\mathbf{W}_{i}\\ \end{bmatrix}\mathbf{Z}_{\mathbf{v}}^{+}+\mathbf{b}_{r}$
Both $\mathcal{L}_{\mathbb{R}\rightarrow\mathbb{C}}$ and $\mathcal{L}_{\mathbb{C}\rightarrow\mathbb{R}}$ can be implemented with simple linear layers, with the former lifting the $d$-dimensional real-valued inputs to $2d$ dimensional queries, keys, and values, and the latter projecting the $2d$-dimensional predictions back to $d$-dimensional real-valued outputs in the original basis. Hence, every step of the Isotropic AFA computation can be performed entirely in the real domain.
@@ 4.2 Algorithm
$\mathbf{W}[i,j]:=\mathbf{W}_{r}[i,j]+i\mathbf{W}_{i}[i,j]=\alpha_{ij}\begin{bmatrix}\cos(\phi_{ij})\\ \sin(\phi_{ij})\end{bmatrix}$,
$\quad\alpha_{ij}\sim\sqrt{\tfrac{2}{d_{\text{in}}+d_{\text{{out}}}}}\,\mathcal{N}(0,1),\;\;\phi_{ij}\sim\mathcal{U}(0,2\pi).$
We use scalar noise parameters $\sigma^{2},\eta^{2}$ and dynamics parameters $\alpha,\bm{\omega}$, where the real part of the eigenvalues is set to $\alpha^{2}$ to ensure stability of our LTI model. Note that the eigenvalues of $\bm{\Omega}$ appear in complex conjugate pairs, as a necessary condition for $\mathbf{A}$ to be real-valued, so we can write the vector of eigenvalues as:
@@
We drop $\big|\lambda_{C}\big|^{2}$ since we can absorb it into $\sigma^{2}$, and we drop the Mahalanobis scale $\nu$, since we can absorb it into $\sigma,\eta,\tau_{s},\mathbf{W}_{Q},\mathbf{W}_{K}$.
We switch from the rational weight $\tfrac{1}{1+r_{ij}^{2}/\nu}$ to the equivalent exponential–log form $\exp\!\big(-\tau_{s}\log(1+r_{ij}^{2}/\nu)\big)$, where $\tau_{s}$ is the usual temperature parameter that controls the sharpness of the weighting, with the aim of improving numerical stability; this reduces exactly to the rational form when $\tau_{s}=1$. (However, this may not be necessary.)
For simplicity, we present the case of shared dynamic parameters $\alpha,\bm{\omega}$ and noise parameters $\sigma,\eta,\gamma$ for the queries/keys and values, so $\bm{\tilde{\Phi}}_{\mathbf{qk}}^{+}=\bm{\tilde{\Phi}}_{\mathbf{v}}^{+}$ and $\bm{\tilde{\Phi}}_{\mathbf{qk}}^{-}=\bm{\tilde{\Phi}}_{\mathbf{v}}^{-}$. We define the latter as separate variables however, since in general they can be different.
To deal with the case of small $\alpha$, we can use the expression from L’Hôpital’s rule and define a masked expression for $\sigma^{2}_{V}$:
$\kappa[\tau]=\frac{1-\mathcal{K}^{2}_{\alpha}[\tau]+\epsilon}{2\alpha^{2}+\epsilon}\mathbf{1}_{\{\alpha^{2}>\epsilon\}}+\tau\mathbf{1}_{\{\alpha^{2}\leq\epsilon\}}$
Our covariance matrix is then defined as:
$\mathbf{V}^{Z}[i,j]=\sigma^{2}_{V}(i-j),\quad\sigma^{2}_{V}(\tau)=\sigma^{2}\kappa[|\tau|]+\eta^{2}\mathcal{K}^{2}_{\alpha}[|\tau|]+\gamma^{2}$
For large context windows, where $e^{\alpha^{2}\Delta t}$ approaches numerical zero for distant tokens, practical implementations may benefit from a small, learned constant floor $\beta^{2}$ added to the decay term, $\mathbf{E}_{\mathbf{qk}}\leftarrow\mathbf{E}_{\mathbf{qk}}+\beta^{2}$.
@@
In what follows, we use the more compact complex notation to represent complex-valued operations, but this should be understood as shorthand for the real-valued computations discussed above.
Input: Input sequence $\mathbf{Z}\in\mathbb{R}^{d\times m}$
Linear layers: $\mathcal{L}_{Q}^{\mathbb{R}\rightarrow\mathbb{C}},\mathcal{L}_{K}^{\mathbb{R}\rightarrow\mathbb{C}},\mathcal{L}_{V}^{\mathbb{R}\rightarrow\mathbb{C}},\mathcal{L}_{P}^{\mathbb{C}\rightarrow\mathbb{R}}$; dynamics/noise parameters: $\alpha,\bm{\omega},\sigma,\eta,\gamma$; time step $\Delta t$; softmax temperature $\tau_{s}$; residual weight $\delta$.
@@
$\big(\operatorname{Re}(\mathbf{Z_{q}}),\operatorname{Im}(\mathbf{Z_{q}})\big)\leftarrow\mathcal{L}_{Q}^{\mathbb{R}\rightarrow\mathbb{C}}\big(\mathbf{Z}\big),$
$\big(\operatorname{Re}(\mathbf{Z_{k}}),\operatorname{Im}(\mathbf{Z_{k}})\big)\leftarrow\mathcal{L}_{K}^{\mathbb{R}\rightarrow\mathbb{C}}\big(\mathbf{Z}\big),$
$\big(\operatorname{Re}(\mathbf{Z_{v}}),\operatorname{Im}(\mathbf{Z_{v}})\big)\leftarrow\mathcal{L}_{V}^{\mathbb{R}\rightarrow\mathbb{C}}\big(\mathbf{Z}\big).$
@@
Covariance kernel: $\mathbf{V}^{Z}[i,j]=\sigma^{2}_{V}(i-j),\quad\sigma^{2}_{V}(\tau)=\sigma^{2}\kappa[|\tau|]+\eta^{2}\mathcal{K}^{2}_{\alpha}[|\tau|]+\gamma^{2}$
@@
$\mathbf{L}=-\log\left(\mathbf{V}^{Z}+\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}+\epsilon\right)\odot\mathbf{M}_{\text{causal}},$
@@
Prediction: $\mathbf{Z}_{\mathbf{v}}^{+}=(e^{\alpha^{2}\Delta t}e^{\bm{\lambda}_{\Omega}\Delta t})\odot\mathbf{\bar{Z}_{v}}$
Output Projection: $\operatorname{Re}(\mathbf{Z_{p}})\leftarrow\mathcal{L}_{P}^{\mathbb{C}\rightarrow\mathbb{R}}\big(\operatorname{Re}(\mathbf{Z}_{\mathbf{v}}^{+}),\operatorname{Im}(\mathbf{Z}_{\mathbf{v}}^{+})\big)$
Return: $\operatorname{Re}(\mathbf{Z_{p}})$ .
@@
    Complex-valued multiplicative positional encodings with learned parameters and scalar decay constants (one per head);
@@
    Pair-wise uncertainty propagation via the closed-form solution of the differential Lyapunov equation;